OpenStudy (thats_kyy):

Solve the system of equations and choose the correct answer from the list of options. d + e = 12 −d + e = −8 Label the ordered pair as (d, e).

OpenStudy (thats_kyy):

- (0,0) - (2,10) - (10,2) - (12,0)

OpenStudy (thats_kyy):

@perl May You please Help Me?

OpenStudy (thats_kyy):

@campbell_st @StudyGurl14

OpenStudy (campbell_st):

if you add the 2 equations you can eliminate d d + e = 12 -d + e = -8 ------------ what do you get...?

OpenStudy (thats_kyy):

20 @campbell_st

OpenStudy (campbell_st):

no not quite d + e = 12 - d + e = -8 -------------- 2e = 4 solve for e. substitute the value for e, into either equation to find d.

OpenStudy (thats_kyy):

Okay

OpenStudy (thats_kyy):

So is the ordered pair (10,2) @campbell_st

OpenStudy (campbell_st):

that's correct

OpenStudy (thats_kyy):

Okay Thank You. @campbell_st